# Deviation Inequalities for Rényi Divergence Estimators via Variational Expression

### v1

@@ 6. Concluding Remarks
We established exponential deviation inequalities for smoothed plug-in estimators and neural estimators of Rényi divergences. The derived inequalities apply to smoothed KL divergence estimate when the unsmoothed distributions are compactly supported or sub-Gaussian, and to smoothed Rényi divergence estimate of order $\alpha\neq 1$ when the distributions are compactly supported. It is worth emphasizing that the dependence of data dimension and distributional parameters are specified in our bounds up to universal constants. However, these dependencies are possibly sub-optimal, and a tighter analysis quantifying their impact maybe warranted in scenarios where parameters are not fixed. For instance, such a scenario naturally arises in physical layer security, where the dimension (codeword length) scales as a function of sample size (codebook size). Also of interest is to determine whether the deviation inequalities are optimal in terms of dependence on $n$, $d$, and other distributional parameters, and otherwise, ascertain the gap from optimality. We hope that our results will be a starting point for further investigations along these lines.
@@ Appendix A Variational expression for Rényi divergences
where the supremum is over all bounded continuous functions $f$. On the other hand, by applying $\log x\leq x-1$ for $x\geq 0$ to last term in (94), we obtain the reverse inequality, thus proving (93). When $\mu\ll\nu$ and $\left(\frac{d\mu}{d\nu}\right)^{\alpha}\in L_{1}(\nu)$, $\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)<\infty$ and the integrals in (93) are finite. Moreover, it can be seen by substitution that the supremum in (93) is achieved by $f_{\alpha}^{\star}=\log\frac{d\mu}{d\nu}+\frac{1-\alpha}{\alpha}\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$. ∎
@@ Appendix B Determination of Prefactor in Theorem 8
The claim in (80) follows by determining the constants appearing in the proof of [75, Theorem 2.7.1]. It is sufficient to consider the case $L=1$ since the claim for general $L$ follows by scaling. Let $\delta=\epsilon^{\frac{1}{\beta}}$, $\left\|\cdot\right\|$ denote the Euclidean norm on $\mathbb{R}^{d}$, and $\{x_{(i)}\}_{i=1}^{m}$, with $m=N(\delta,\mathbb{B}_{d}(r),\left\|\cdot\right\|)$, denote the set of points in the interior of $\mathbb{B}_{d}(r)$ such that balls of radius $\delta$ centered at these points cover $\mathbb{B}_{d}(r)$. We will use the fact that $m\leq\left(1+\frac{2r}{\delta}\right)^{d}$ for $\delta<r$ and $m=1$ for $\delta\geq r$, which implies that $m\leq 1\vee\left(\frac{3r}{\delta}\right)^{d}$. Recall the differential operator $D^{k}$ for a $d$-dimensional vector $k$ with non-negative integers as elements (see Definition 1). Set $\frac{x^{k}}{k!}:=\prod_{i=1}^{d}\frac{x_{i}^{k_{i}}}{k_{i}!}$. For $k$ with $\left|k\right|\leq\underaccent{\bar}{\beta}$, and for each $f$, consider the vector
@@
For any $x\in\mathbb{B}_{d}(r)$, there exists an $x_{(i)}$ for some $1\leq i\leq m$ such that $\left\|x-x_{(i)}\right\|\leq\delta$. By multivariate Taylor’s theorem with remainder estimates given in [23, Equation 8], for $f,g\in C_{1}^{\beta}\big{(}\mathbb{B}_{d}(r)\big{)}$ and $x\in\mathrm{int}\big{(}\mathbb{B}_{d}(r)\big{)}$, we have
@@
$R_{\beta}(f,x_{i},x-x_{(i)})=\underaccent{\bar}{\beta}\sum_{|k|=\underaccent{\bar}{\beta}}\frac{\big{(}x-x_{(i)}\big{)}^{k}}{k!}\int_{0}^{1}(1-t)^{\underaccent{\bar}{\beta}-1}\Big{(}D^{k}f\big{(}(1-t)x_{(i)}+tx\big{)}-D^{k}f\big{(}x_{(i)}\big{)}\Big{)}dt.$ (96)
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left|D^{k}f\big{(}x_{(j)}\big{)}-\sum_{\left|l\right|\leq\underaccent{\bar}{\beta}-|k|}B^{k+l}f\big{(}x_{(i)}\big{)}\frac{\big{(}x_{(j)}-x_{(i)}\big{)}^{l}}{l!}\right|+\left|D^{k}g\big{(}x_{(j)}\big{)}-\sum_{\left|l\right|\leq\underaccent{\bar}{\beta}-|k|}B^{k+l}g\big{(}x_{(i)}\big{)}\frac{\big{(}x_{(j)}-x_{(i)}\big{)}^{l}}{l!}\right|$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{\left|l\right|\leq\underaccent{\bar}{\beta}-|k|}\left|D^{k+l}f\big{(}x_{(i)}\big{)}-B^{k+l}f\big{(}x_{(i)}\big{)}\right|\frac{\left|\big{(}x_{(j)}-x_{(i)}\big{)}^{l}\right|}{l!}+\left|D^{k+l}g\big{(}x_{(i)}\big{)}-B^{k+l}g\big{(}x_{(i)}\big{)}\right|\frac{\left|\big{(}x_{(j)}-x_{(i)}\big{)}^{l}\right|}{l!}$
@@
Given (80), the upper bound in (81) follows via similar steps given in [75, Corollary 2.7.2], which deals with bracketing entropy in place of covering entropy.
@@ C.2. Proof of Lemma 6
We will follow the proof given in [75, Lemma 2.2.2.], by avoiding arguments that require convexity of $\psi$. For $\psi_{q}(x)=e^{x^{q}}-1$, let $x^{\star}_{q}=\big{(}\log(3/2)\big{)}^{1/q}$. Observe that for $x,y\geq x^{\star}_{q}$ and $c_{q}^{\prime}\geq\big{(}2/\log(3/2)\big{)}^{1/q}$, we have $\psi_{q}(x)\psi_{q}(y)\leq\psi_{q}(c_{q}^{\prime}xy)$. Consequently, for any $y\geq x^{\star}_{q}$ and $x\geq x^{\star}_{q}y$,

### v3

@@ 6. Concluding Remarks
We established exponential deviation inequalities for smoothed plug-in estimators and neural estimators of Rényi divergences. The derived inequalities apply to smoothed KL divergence estimate when the unsmoothed distributions are compactly supported or sub-Gaussian, and to smoothed Rényi divergence estimate of order $\alpha\neq 1$ when the unsmoothed distributions are compactly supported. It is worth emphasizing that the dependence of data dimension and distributional parameters are specified in our bounds up to universal constants. However, these dependencies are possibly sub-optimal, and a tighter analysis quantifying their impact maybe warranted in scenarios where parameters are not fixed. For instance, such a scenario naturally arises in physical-layer security, where the dimension (codeword length) scales as a function of sample size (codebook size). A plausible approach in this regard is to find a suitable class of smooth functions with a better covering entropy to which the supremum in the variational form of Rényi divergences can be restricted to. Another interesting direction to pursue is deviation inequalities for kernel density estimators via the variational approach. We hope that our results will be a starting point for further investigations along these lines.
@@ Appendix A Variational Expression for Rényi Divergences
where the supremum is over all bounded continuous functions $f$. On the other hand, by applying $\log x\leq x-1$ for $x\geq 0$ to the last term in (89), we obtain the reverse inequality, thus proving (88). When $\mu\ll\nu$ and $\left(\frac{d\mu}{d\nu}\right)^{\alpha}\in L_{1}(\nu)$, $\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)<\infty$ and the integrals in (88) are finite. Moreover, it can be seen by substitution that the supremum in (88) is achieved by $f_{\alpha}^{\star}=\log\frac{d\mu}{d\nu}+\frac{1-\alpha}{\alpha}\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$. ∎
@@ Appendix B Determination of Prefactor in Theorem 8
The claim in (74) will follow by determining the constants appearing in the proof of [83, Theorem 2.7.1]. It is sufficient to consider the case $L=1$ since the claim for general $L$ follows by scaling. Let $\delta=\epsilon^{\frac{1}{\beta}}$, $\left\|\cdot\right\|$ denote the Euclidean norm on $\mathbb{R}^{d}$, and $\{x_{(i)}\}_{i=1}^{m}$, with $m=N(\delta,\mathbb{B}_{d}(r),\left\|\cdot\right\|)$, denote the set of points in the interior of $\mathbb{B}_{d}(r)$ such that balls of radius $\delta$ centered at these points cover $\mathbb{B}_{d}(r)$. We will use the fact that $m\leq\left(1+\frac{2r}{\delta}\right)^{d}$ for $\delta<r$ and $m=1$ for $\delta\geq r$, which implies that $m\leq 1\vee\left(\frac{3r}{\delta}\right)^{d}$. Recall the differential operator $D^{k}$ for a $d$-dimensional vector $k$ with non-negative integers as elements (see Definition 1). Set $\frac{x^{k}}{k!}:=\prod_{i=1}^{d}\frac{x_{i}^{k_{i}}}{k_{i}!}$. For $k$ with $\left|k\right|\leq\underaccent{\bar}{\beta}$, and for each $f$, consider the vector
@@
For any $x\in\mathbb{B}_{d}(r)$, there exists an $x_{(i)}$ for some $1\leq i\leq m$ such that $\left\|x-x_{(i)}\right\|\leq\delta$. By multivariate Taylor’s theorem with remainder estimates given in [28, Equation 8], for $f,g\in C_{1}^{\beta}\big{(}\mathbb{B}_{d}(r)\big{)}$ and $x\in\mathrm{int}\big{(}\mathbb{B}_{d}(r)\big{)}$, we have
@@
$R_{\beta}(f,x_{(i)},x-x_{(i)})=\underaccent{\bar}{\beta}\sum_{|k|=\underaccent{\bar}{\beta}}\frac{\big{(}x-x_{(i)}\big{)}^{k}}{k!}\int_{0}^{1}(1-t)^{\underaccent{\bar}{\beta}-1}\Big{(}D^{k}f\big{(}(1-t)x_{(i)}+tx\big{)}-D^{k}f\big{(}x_{(i)}\big{)}\Big{)}dt.$ (91)
@@
[…]
@@
Given (74), the upper bound in (75) follows via similar steps given in [83, Corollary 2.7.2], which deals with bracketing entropy in place of covering entropy.
@@ C.2. Proof of Lemma 6
We will follow the proof given in [83, Lemma 2.2.2.], by avoiding arguments that require convexity of $\psi$. For $\psi_{q}(x)=e^{x^{q}}-1$, let $x^{\star}_{q}=\big{(}\log(3/2)\big{)}^{1/q}$. Observe that for $x,y\geq x^{\star}_{q}$ and $c_{q}^{\prime}\geq\big{(}2/\log(3/2)\big{)}^{1/q}$, we have $\psi_{q}(x)\psi_{q}(y)\leq\psi_{q}(c_{q}^{\prime}xy)$. Consequently, for any $y\geq x^{\star}_{q}$ and $x\geq x^{\star}_{q}y$,